# Spurious Correlations in High Dimensional Regression: The Roles of Regularization, Simplicity Bias and Over-Parameterization

### v1

@@ 1 Introduction
Throughout the paper, the theoretical results are supported by numerical experiments on Gaussian data, Color-MNIST, and CIFAR-10, which validates our analysis even in settings not strictly following the modeling choices.
@@ 2 Related work
In contrast with linear regression where the number of parameters equals the input dimension, random features models [44] capture the effects of over-parameterization, as the number of parameters is independently of $d$ and $n$. [32] have characterized the test loss of random features, showing that it displays a double descent [5]. Furthermore, the RF model has been used to understand a wide family of phenomena such as feature learning [3, 13, 34], robustness under adversarial attacks [14, 7, 20], and distribution shift [54, 28]. The equivalence between an over-parameterized RF model and a regularized linear one has also been studied in detail [17, 18, 24, 36]. However, existing rigorous results show the equivalence at the level of training and test error. In contrast, we are interested in the covariance defined in (3.3) and, for this reason, we prove an equivalence at the level of the predictor (Theorem 2).
@@ 3 Preliminaries
Given a vector $v$, we denote by $\left\|v\right\|_{2}$ its Euclidean norm. Given a matrix $A$, we denote by $\mathop{\rm tr}\nolimits(A)$ and $\left\lVert A\right\rVert_{\textup{op}}$ its trace and operator (spectral) norm. Given a symmetric matrix $A$, we denote by $\lambda_{\rm min}\left(A\right)$ ($\lambda_{\rm max}\left(A\right)$) its smallest (largest) eigenvalue. All complexity notations $\Omega(\cdot)$, $\mathcal{O}(\cdot)$, $\omega(\cdot)$, $o(\cdot)$ and $\Theta(\cdot)$ are understood for large data size $n$, input dimension $d$, and number of parameters $p$. We indicate with $C,c>0$ numerical constants, independent of $n$, $d$, $p$, whose value may change from line to line.
@@
We consider supervised learning with $n$ training samples $\{(z_{1},g_{1}),\ldots,(z_{n},g_{n})\}$ and labels defined by a (not necessarily deterministic) function of the inputs $g_{i}=f^{*}(z_{i})$, where $z_{i}\in\mathbb{R}^{2d}$ denotes the $i$-th training input and $g_{i}\in\mathbb{R}$ the corresponding label. Input samples are composed by two distinct parts (or features), i.e., $z_{i}^{\top}=[x_{i}^{\top},y_{i}^{\top}]$, with $x_{i},y_{i}\in\mathbb{R}^{d}$, and they are sampled i.i.d. from the distribution $\mathcal{P}_{XY}$. We further denote with $\mathcal{P}_{X}$ ($\mathcal{P}_{Y}$) the marginal distribution of the $x_{i}$-s ($y_{i}$-s). The features $x$ and $y$ have the same dimension $d$ to ease the presentation.
We focus on the setting where the labels $g_{i}$ depend only on $x_{i}$, i.e., $g_{i}=f^{*}(z_{i})=f^{*}_{x}(x_{i})$ for some (not necessarily deterministic) function $f^{*}_{x}$. Hence, $y_{i}$ is independent from $g_{i}$, after conditioning on $x_{i}$. We highlight that the independence between $y_{i}$ and $g_{i}$ is conditional to $x_{i}$, as the covariance between $y_{i}$ and $x_{i}$ is in general non-zero. We refer to $y_{i}$ as the spurious feature of the $i$-th sample, and to $x_{i}$ as its core feature. As an example, $x_{i}$ may represent the main object in an image and $y_{i}$ the (not necessarily independent) background, see Figure 1.
@@
for some regularization term $\lambda\geq 0$, where $\ell$ is a loss function [^1] In general, existence and uniqueness of $\hat{\theta}$ depend on the choice of the model $f(\theta,z)$, the loss function $\ell$ and the regularization term $\lambda$. For the purposes of our work, we will precisely define $\hat{\theta}$ for linear regression (Section 4) and for random features (Section 6).. We define the (in-distribution) test loss associated to the model $f(\hat{\theta},\cdot)$ as
@@
This implies that an increase in $\mathcal{C}(\hat{\theta})$ hurts the performance of the model when core and spurious features are sampled independently (and, thus, the model is tested out-of-distribution). The proof of (3.4) is in Appendix D.
@@ 4 Precise analysis for linear regression
Let $\lambda=0$ and $Z^{\top}Z\in\mathbb{R}^{2d\times 2d}$ be invertible [^2] Under Assumption 1, this holds with probability 1 for $n\geq 2d$.. Let $\mathcal{C}(\hat{\theta}_{\textup{LR}}(0))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(0))$. Then, we have that
@@
In words, $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(0))$ does not learn any spurious correlation between the spurious feature $y$ and the label $g$. This is also clear from Figure 2, where we report in red the value of $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$, which approaches 0 as $\lambda$ becomes small.
@@
Let Assumption 1 hold, $n=\Theta(d)$ and $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda))$ for $\lambda>0$. Denote by $P_{y}\in\mathbb{R}^{2d\times 2d}$ the projector on the last $d$ elements of the canonical basis in $\mathbb{R}^{2d}$, and set
@@
Figure 2: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda)$ (red) as a function of the regularization term $\lambda$ for two values of the number of samples $n$. Left: synthetic Gaussian dataset, with $d=400$ (additional details in Appendix E); right: binary Color-MNIST dataset with correlation $\sqrt{1-\alpha^{2}}=0.25$ between color and digit (see Figure 1).
In words, Theorem 1 guarantees that $|\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))-\mathcal{C}^{\Sigma}(\lambda)|=o(1)$ with high probability (e.g., setting $t=d^{-1/5}$). Thus, for large $d,n$, we can estimate $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ via the deterministic quantity $\mathcal{C}^{\Sigma}(\lambda)$, which depends on the true parameter $\theta^{*}$, the covariance of the data $\Sigma$, and the regularization $\lambda$ via the parameter $\tau(\lambda)$ introduced in (4.9).
Note that, since $\hat{\theta}_{\textup{LR}}(\lambda)$ is given by (4.4), when $\lambda>0$ it cannot be decomposed as $\theta^{*}+\left(Z^{\top}Z\right)^{-1}Z^{\top}\mathcal{E}$ (as in the proof of Proposition 4.1 for $\lambda=0$). Thus, we rely on the non-asymptotic characterization of $\hat{\theta}_{\textup{LR}}(\lambda)$ recently provided by [19]. In particular, their analysis focuses on the proportional regime $n=\Theta(d)$, and it allows to provide concentration bounds on a certain family of low-dimensional functions of $\hat{\theta}_{\textup{LR}}(\lambda)$, which includes $\mathcal{C}$ as defined in (3.3). The details are in Appendix B.
@@ 5 Roles of regularization and simplicity bias
Figure 3: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (red) as a function of $\lambda_{\rm max}\left(\Sigma_{yy}\right)$ (left) and $\lambda_{\rm min}\left(S_{x}^{\Sigma}\right)$ (right) on a synthetic Gaussian dataset, for both linear regression and random features, with $\lambda=1$ (additional details in Appendix E).
@@
Let Assumption 1 hold, $n=\Theta(d)$ and $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the in-distribution test loss of the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda))$ for $\lambda>0$. Set
@@
Figure 4: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda)$ (red) as a function of $\lambda_{\rm max}\left(\Sigma_{yy}\right)/\mathop{\rm tr}\nolimits(\Sigma_{yy})$ on a CIFAR-10 dataset for different levels of whitening (details on the whitening process in Appendix E). We restrict to the classes “boat” and “truck” ($n=10000$) and consider both linear regression and random features, with $\lambda=1$.
@@
The same conclusion does not hold for the out-of-distribution test loss, where the features $x$ and $y$ are sampled independently: the lower bound in (3.4) increases with $\mathcal{C}$. Figure 6 in Appendix D provides an additional numerical validation of this statement.
@@ 6 Role of over-parameterization
Figure 5: Test loss $\mathcal{L}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda)$ (red) as a function of $\lambda$. Left: 2-layer fully connected ReLU network, trained on the binary color(C)-MNIST and CIFAR-10 (boats and trucks). Right: RF model with $\tanh$ and $\phi_{1}=h_{1}+0.1\,h_{3}$ activation. Implementation details are in Appendix E.
@@
$f_{\textup{RF}}(z,\theta)=\phi(Vz)^{\top}\theta,$ (6.1)
@@
Let Assumptions 2, 3, and 4 hold, $n=\Theta(d)$, and $z\in\mathbb{R}^{2d}$ be sampled from a distribution satisfying Assumption 4, not necessarily with the same covariance as $\mathcal{P}_{XY}$, independent from everything else. Let $f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))$ be the RF model defined in (6.1) with $\hat{\theta}_{\textup{RF}}(\lambda)$ given by (6.3), and $f_{\textup{LR}}(z,\hat{\theta}_{\textup{LR}}(\tilde{\lambda}))$ be the linear regression model defined in (4.1) with $\hat{\theta}_{\textup{LR}}(\tilde{\lambda})$ given by (4.4). Then, for $\lambda\geq 0$,
$\left|f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))-f_{\textup{LR}}(z,\hat{\theta}_{\textup{LR}}(\tilde{\lambda}))\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right)=o(1),$ (6.4)
@@ Appendix B Proofs for linear regression
$\hat{\theta}_{\textup{LR}}(0)=\left(Z^{\top}Z\right)^{-1}Z^{\top}\left(Z\theta^{*}+\epsilon\right)=\theta^{*}+\left(Z^{\top}Z\right)^{-1}Z^{\top}\mathcal{E}.$ (B.2)
@@
where $\rho$ is a standard Gaussian vector in $\mathbb{R}^{2d}$. In the equation above, $\gamma>0$ is implicitly defined via
$\frac{n\gamma^{2}}{2d}=\sigma^{2}+\mathbb{E}_{\rho}\left[\left\|\Sigma^{1/2}\left(\hat{\theta}^{\rho}-\theta^{*}\right)\right\|_{2}^{2}\right]$ (B.10)
@@
$=\theta^{\top}\mathbb{E}_{[x^{\top},y^{\top}]^{\top}\sim P_{XY}}\left[[\mathbf{0}^{\top},y^{\top}]^{\top}[x^{\top},\mathbf{0}^{\top}]^{\top}\right]\theta^{*}$
$=\theta^{\top}P_{y}\mathbb{E}_{[x^{\top},y^{\top}]^{\top}\sim P_{XY}}\left[[x^{\top},y^{\top}]^{\top}[x^{\top},y^{\top}]^{\top}\right]\theta^{*}$
@@
where the third line holds since $\tilde{x}$ has 0 mean and is independent with $x$ and $y$, and by definition of $\theta^{*}_{x}$, and the fourth line holds because $P_{y}[x^{\top},y^{\top}]^{\top}=[\mathbf{0}^{\top},y^{\top}]^{\top}$ and because the last $d$ entries of $\theta^{*}$ are 0 (i.e., $P_{y}\theta^{*}=0$). Thus, since we have that $\left\|P_{y}\Sigma\theta^{*}\right\|_{2}\leq\left\lVert P_{y}\right\rVert_{\textup{op}}\left\lVert\Sigma\right\rVert_{\textup{op}}\left\|\theta^{*}\right\|_{2}\leq\left\lVert\Sigma\right\rVert_{\textup{op}}$ because of Assumption 1, we have that $\mathcal{C}(\cdot):\mathbb{R}^{2d}\to\mathbb{R}$ is a $\left\lVert\Sigma\right\rVert_{\textup{op}}$-Lipschitz function.
Now, since $\mathcal{P}_{XY}$ is multivariate Gaussian, Theorem 2.3 of [19] gives that, for any 1-Lipschitz function $\varphi:\mathbb{R}^{2d}\to\mathbb{R}$, and any $t\in(0,1/2)$,
@@
$\leq\tau\left\|\left(\Sigma_{xx}+\tau I\right)^{-1/2}\theta_{x}^{*}\right\|_{2}\left\|\left(\Sigma_{xx}+\tau I\right)^{1/2}\theta^{*}_{x}\right\|_{2}$
$\qquad\left\lVert\left(\Sigma_{xx}+\tau I\right)^{-1/2}\Sigma_{xy}\left(S_{x}^{\Sigma+\tau I}\right)^{-1}\Sigma_{yx}\left(\Sigma_{xx}+\tau I\right)^{-1/2}\right\rVert_{\textup{op}}$
$\leq\tau\,\frac{1}{\sqrt{\lambda_{\rm min}\left(\Sigma_{xx}\right)+\tau}}\,\sqrt{{\theta^{*}_{x}}^{\top}\Sigma_{xx}\theta^{*}_{x}+\tau}$
$\qquad\left\lVert\left(\Sigma_{xx}+\tau I\right)^{-1/2}\Sigma_{xy}\left(S_{x}^{\Sigma+\tau I}\right)^{-1}\Sigma_{yx}\left(\Sigma_{xx}+\tau I\right)^{-1/2}\right\rVert_{\textup{op}}$

### v2

@@ 1 Introduction
Throughout the paper, the theoretical results are supported by numerical experiments on synthetic Gaussian data, Color-MNIST, and CIFAR-10, which validates our analysis even in settings not strictly following the modeling choices. Our code is publicly available at the GitHub repository simone-bombari/spurious-correlations.
@@ 2 Related work
In contrast with linear regression where the number of parameters equals the input dimension, random features models [46] capture the effects of over-parameterization, as the number of parameters is independent of $d$ and $n$. [34] have characterized the test loss of random features, showing that it displays a double descent [5]. Furthermore, the RF model has been used to understand a wide family of phenomena such as feature learning [3, 13, 36], robustness under adversarial attacks [14, 7, 20], and distribution shift [56, 28]. This model has been also considered in the setting where the data has two dependent components [31], although it was used to study the training and generalization error when only partial information is available. The equivalence between an over-parameterized RF model and a regularized linear one has also been studied in detail [17, 18, 24, 38]. However, existing rigorous results show the equivalence at the level of training and test error. In contrast, we are interested in the covariance defined in (3.3) and, for this reason, we prove an equivalence at the level of the predictor (Theorem 2).
@@ 3 Preliminaries
Given a vector $v$, we denote by $\left\|v\right\|_{2}$ its Euclidean norm. Given a matrix $A$, we denote by $\mathop{\rm tr}\nolimits(A)$ and $\left\lVert A\right\rVert_{\textup{op}}$ its trace and operator (spectral) norm. Given a symmetric matrix $A$, we denote by $\lambda_{\rm min}\left(A\right)$ ($\lambda_{\rm max}\left(A\right)$) its smallest (largest) eigenvalue. We denote by $\lambda$ (without a sub-script) the $\ell_{2}$ regularization term in ridge regression. All complexity notations $\Omega(\cdot)$, $\mathcal{O}(\cdot)$, $\omega(\cdot)$, $o(\cdot)$ and $\Theta(\cdot)$ are understood for large data size $n$, input dimension $d$, and number of parameters $p$. We indicate with $C,c>0$ numerical constants, independent of $n$, $d$, $p$, whose value may change from line to line.
@@
We consider supervised learning with $n$ training samples $\{(z_{1},g_{1}),\ldots,(z_{n},g_{n})\}$ and labels defined by a (not necessarily deterministic) function of the inputs $g_{i}=f^{*}(z_{i})$, where $z_{i}\in\mathbb{R}^{2d}$ denotes the $i$-th training input and $g_{i}\in\mathbb{R}$ the corresponding label. Input samples are composed by two distinct parts (or features), i.e., $z_{i}^{\top}=[x_{i}^{\top},y_{i}^{\top}]$, with $x_{i},y_{i}\in\mathbb{R}^{d}$, and they are sampled i.i.d. from the distribution $\mathcal{P}_{XY}$. We further denote with $\mathcal{P}_{X}$ ($\mathcal{P}_{Y}$) the marginal distribution of the $x_{i}$-s ($y_{i}$-s). The features $x$ and $y$ have the same dimension $d$ to ease the presentation. We opted to focus on the setting $z=[x,y]$ due to its connection with the motivating example of images with background, but the analysis could be extended to other cases, such as $z=x+y$, briefly discussed in Appendix D.
We focus on the setting where the labels $g_{i}$ depend only on $x_{i}$, i.e., $g_{i}=f^{*}(z_{i})=f^{*}_{x}(x_{i})$ for some (not necessarily deterministic) function $f^{*}_{x}$. Hence, $y_{i}$ is independent from $g_{i}$, after conditioning on $x_{i}$. We highlight that the independence between $y_{i}$ and $g_{i}$ is conditional on $x_{i}$, and that the covariance between $y_{i}$ and $x_{i}$ is in general non-zero. We refer to $y_{i}$ as the spurious feature of the $i$-th sample, and to $x_{i}$ as its core feature. As an example, $x_{i}$ may represent the main object in an image and $y_{i}$ the (not necessarily independent) background, see Figure 1.
@@
for some regularization term $\lambda\geq 0$, where $\ell$ is a loss function [^1] In general, existence and uniqueness of $\hat{\theta}$ depend on the choice of the model $f(\theta,z)$, the loss function $\ell$ and the regularization term $\lambda$. For the purposes of our work, we will precisely define $\hat{\theta}$ for linear regression (Section 4) and for random features (Section 6).. We define the test loss associated to the model $f(\hat{\theta},\cdot)$ as
@@
We note that this test loss is in distribution.
@@
This implies that an increase in $\mathcal{C}(\hat{\theta})$ hurts the performance of the model when core and spurious features are sampled independently (and, thus, the model is tested out-of-distribution). The proof of (3.4) is in Appendix E.
@@ 4 Precise analysis for linear regression
Let $\lambda=0$ and $Z^{\top}Z\in\mathbb{R}^{2d\times 2d}$ be invertible [^2] Under Assumption 1, this holds with probability 1 for $n\geq 2d$.. Let $\mathcal{C}(\hat{\theta}_{\textup{LR}}(0))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(0),\cdot)$. Then, we have that
@@
In words, $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(0),\cdot)$ does not learn any spurious correlation between the spurious feature $y$ and the label $g$. This is also clear from Figure 2, where we report in red the value of $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$, which approaches 0 as $\lambda$ becomes small.
@@
Let Assumption 1 hold, $n=\Theta(d)$ and $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda),\cdot)$ for $\lambda>0$. Denote by $P_{y}\in\mathbb{R}^{2d\times 2d}$ the projector on the last $d$ elements of the canonical basis in $\mathbb{R}^{2d}$, and set
@@
Figure 2: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda)$ (red) as a function of the regularization term $\lambda$ for two values of the number of samples $n$. Left: synthetic Gaussian dataset, with $d=400$ (additional details in Appendix F); right: binary Color-MNIST dataset with correlation $\sqrt{1-\alpha^{2}}=0.25$ between color and digit (see Figure 1).
In words, Theorem 1 guarantees that $|\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))-\mathcal{C}^{\Sigma}(\lambda)|=o(1)$ with high probability (e.g., setting $t=d^{-1/5}$). Thus, for large $d,n$, we can theoretically analyze $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ via the deterministic quantity $\mathcal{C}^{\Sigma}(\lambda)$, which, as highlighted by (4.8), depends on $\theta^{*}$, the covariance of the data $\Sigma$, and the regularization $\lambda$ via the parameter $\tau(\lambda)$ introduced in (4.9). We further analyze the object $\mathcal{C}^{\Sigma}(\lambda)$ in Section 5 that immediately follows.
Note that, since $\hat{\theta}_{\textup{LR}}(\lambda)$ is given by (4.4), when $\lambda>0$ it cannot be decomposed as $\theta^{*}+\left(Z^{\top}Z\right)^{-1}Z^{\top}\mathcal{E}$ (as in the proof of Proposition 4.1 for $\lambda=0$). Thus, we rely on the non-asymptotic characterization of $\hat{\theta}_{\textup{LR}}(\lambda)$ recently provided by [19]. In particular, their analysis focuses on the proportional regime $n=\Theta(d)$ (as opposed to the regime $n=\omega(d)$, considered in (4.6)), and it allows to provide concentration bounds on a certain family of low-dimensional functions of $\hat{\theta}_{\textup{LR}}(\lambda)$, which includes $\mathcal{C}$ as defined in (3.3). The details are in Appendix B.
@@ 5 Roles of regularization and simplicity bias
Figure 3: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (red) as a function of $\lambda_{\rm max}\left(\Sigma_{yy}\right)$ (left) and $\lambda_{\rm min}\left(S_{x}^{\Sigma}\right)$ (right) on a synthetic Gaussian dataset, for both linear regression and random features, with $\lambda=1$ (additional details in Appendix F).
@@
We remark that our results concern the parameter $\hat{\theta}$ as defined in (3.1), which can be interpreted as the convergence point of an optimization algorithm such as gradient descent. This perspective differs from the prior work of [42, 44], which focus on how spurious correlations evolve during training. On the other hand, in linear regression, solving the gradient flow equation $\textup{d}\theta=-\nabla_{\theta}\mathcal{L}_{\lambda}(\theta)\textup{d}t$, where $\mathcal{L}_{\lambda}(\theta)$ is defined as the argument of the $\arg\mathop{\rm min}\nolimits$ in (3.1), gives
$\theta(t)=\left(1-\exp\left(-2\left(X^{\top}X/n+\lambda I\right)t\right)\right)\hat{\theta}.$ (5.6)
Thus, the components of $\hat{\theta}$ aligned with the top eigen-spaces of $X^{\top}X$ converge earlier than the others. Hence, from a dynamical point of view, if $X^{\top}X\sim n\Sigma$, our results suggest that spurious features are learned faster the easier they are.
@@
Figure 4: Test loss $\mathcal{L}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{LR}/\textup{RF}}(\lambda)$ (red) as a function of $\lambda_{\rm max}\left(\Sigma_{yy}\right)/\mathop{\rm tr}\nolimits(\Sigma_{yy})$ on a CIFAR-10 dataset for different levels of whitening (details on the whitening process in Appendix F). We restrict to the classes “boat” and “truck” ($n=10000$) and consider both linear regression and random features, with $\lambda=1$.
Figure 5: Test loss $\mathcal{L}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda)$ (red) as a function of $\lambda$. Left: 2-layer fully connected ReLU network, trained on the binary color(C)-MNIST and CIFAR-10 (boats and trucks). Right: RF model with $\tanh$ and $\phi_{1}=h_{1}+0.1\,h_{3}$ activation. Implementation details are in Appendix F.
@@
Let Assumption 1 hold, $n=\Theta(d)$ and $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the in-distribution test loss of the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda),\cdot)$ for $\lambda>0$. Set
@@
The same conclusion does not hold for the out-of-distribution test loss, where the features $x$ and $y$ are sampled independently: the lower bound in (3.4) increases with $\mathcal{C}$. Figure 6 in Appendix E provides an additional numerical validation of this statement.
@@ 6 Role of Over-parameterization
$f_{\textup{RF}}(\theta,z)=\phi(Vz)^{\top}\theta,$ (6.1)
@@
Let Assumptions 2, 3, and 4 hold, $n=\Theta(d)$, and $z\in\mathbb{R}^{2d}$ be sampled from a distribution satisfying Assumption 4, not necessarily with the same covariance as $\mathcal{P}_{XY}$, independent from everything else. Let $f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)$ be the RF model defined in (6.1) with $\hat{\theta}_{\textup{RF}}(\lambda)$ given by (6.3), and $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\tilde{\lambda}),z)$ be the linear regression model defined in (4.1) with $\hat{\theta}_{\textup{LR}}(\tilde{\lambda})$ given by (4.4). Then, for $\lambda\geq 0$,
$\left|f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)-f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\tilde{\lambda}),z)\right|$ (6.4)
$\hskip 40.00006pt=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right)=o(1),$
@@ Appendix B Proofs for Linear Regression
$\hat{\theta}_{\textup{LR}}(0)=\left(Z^{\top}Z\right)^{-1}Z^{\top}\left(Z\theta^{*}+\mathcal{E}\right)=\theta^{*}+\left(Z^{\top}Z\right)^{-1}Z^{\top}\mathcal{E}.$ (B.2)
@@
where $\rho$ is a standard Gaussian vector in $\mathbb{R}^{2d}$ (the Gaussian sequence model is defined in Equation (1.5) in [19], via the different notation ${\hat{\mu}}^{\textup{seq}}_{(\Sigma,\mu_{0})}$, and our Gaussian vector $\rho$ is denoted as $g$ in the same equation). In the equation above, $\gamma>0$ is implicitly defined via
$\frac{n\gamma^{2}}{2d}=\sigma^{2}+\mathbb{E}_{\rho}\left[\left\|\Sigma^{1/2}\left(\hat{\theta}^{\rho}-\theta^{*}\right)\right\|_{2}^{2}\right].$ (B.10)
@@
$=\theta^{\top}\mathbb{E}_{[x^{\top},y^{\top}]^{\top}\sim P_{XY}}\left[[\mathbf{0}^{\top},y^{\top}]^{\top}[x^{\top},\mathbf{0}^{\top}]\right]\theta^{*}$
$=\theta^{\top}P_{y}\mathbb{E}_{[x^{\top},y^{\top}]^{\top}\sim P_{XY}}\left[[x^{\top},y^{\top}]^{\top}[x^{\top},y^{\top}]\right]\theta^{*}$
@@
where the third line holds since $\tilde{x}$ has 0 mean and is independent from $x$ and $y$, and by definition of $\theta^{*}_{x}$, and the fourth line holds because $P_{y}[x^{\top},y^{\top}]^{\top}=[\mathbf{0}^{\top},y^{\top}]^{\top}$ and because the last $d$ entries of $\theta^{*}$ are 0 (i.e., $P_{y}\theta^{*}=0$). Thus, since we have that $\left\|P_{y}\Sigma\theta^{*}\right\|_{2}\leq\left\lVert P_{y}\right\rVert_{\textup{op}}\left\lVert\Sigma\right\rVert_{\textup{op}}\left\|\theta^{*}\right\|_{2}\leq\left\lVert\Sigma\right\rVert_{\textup{op}}$ because of Assumption 1 (and since $\left\|\theta^{*}\right\|\leq 1$), we have that $\mathcal{C}(\cdot):\mathbb{R}^{2d}\to\mathbb{R}$ is a $\left\lVert\Sigma\right\rVert_{\textup{op}}$-Lipschitz function.
Now, since $\mathcal{P}_{XY}$ is multivariate Gaussian, Theorem 2.3 of [19] gives that, for any 1-Lipschitz function $\varphi:\mathbb{R}^{2d}\to\mathbb{R}$ (denoted as $g$ in [19]), and any $t\in(0,1/2)$,
@@
$\leq\tau\left\|\left(\Sigma_{xx}+\tau I\right)^{-1/2}\theta_{x}^{*}\right\|_{2}\left\|\left(\Sigma_{xx}+\tau I\right)^{1/2}\theta^{*}_{x}\right\|_{2}\left\lVert\left(\Sigma_{xx}+\tau I\right)^{-1/2}\Sigma_{xy}\left(S_{x}^{\Sigma+\tau I}\right)^{-1}\Sigma_{yx}\left(\Sigma_{xx}+\tau I\right)^{-1/2}\right\rVert_{\textup{op}}$
$\leq\tau\,\frac{1}{\sqrt{\lambda_{\rm min}\left(\Sigma_{xx}\right)+\tau}}\,\sqrt{{\theta^{*}_{x}}^{\top}\Sigma_{xx}\theta^{*}_{x}+\tau}\,\left\lVert\left(\Sigma_{xx}+\tau I\right)^{-1/2}\Sigma_{xy}\left(S_{x}^{\Sigma+\tau I}\right)^{-1}\Sigma_{yx}\left(\Sigma_{xx}+\tau I\right)^{-1/2}\right\rVert_{\textup{op}}$